OpenStudy (anonymous):

Calculate the slope of the two points: (-2,-1) and (3,4)

OpenStudy (anonymous):

\[m = \frac{y_{1} - y_{2}}{x_{1} - x_{2}} = \frac{(4) - (-1)}{(3) - (-2)} = \frac{4 + 1}{3 + 2} = \frac??\]

OpenStudy (anonymous):

5/5 ?

OpenStudy (anonymous):

Which can simplify to what?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

There you go :)

OpenStudy (anonymous):

Thank youu